OpenStudy (anonymous):

6/18 what is it in decimal form

OpenStudy (unklerhaukus):

6/18=1/3=0.333...

OpenStudy (anonymous):

simplify 6/18 is the same as 1/3 which is 0.333....

OpenStudy (karatechopper):

0.33333333333333..........

OpenStudy (unklerhaukus):